Is the following an equation of state? (P+ TỚI) (ỹ – b) = RT
Answer #1

+ Cpt ) CV-b) = RT -- 0 as where a cand b are called van der Waals constant simplify elu ☺ (p+ 2 2 2 ) ( - b) = RT Pv - b + cBester - 2004 c multiply) V 3 av - Derb 13 a TV-25] Codeling-bath sa CV p) a a[v-ghob th] 13 -a [v-3b+b = a [v-v+b] 13 Pe = a- 3 3XT9h2 RT - 3b-b CV = 3 b) 22 - 6 276²12 27 br Te = sar! ga 27 bR => The constant in the reen dey wants egn can be dveelvkeчe v = 3, pul is ci* (1) CP + 2 2 2 ) ( 3h - b) = RT 1 - Т. (Put value of a 1. . 2 b. трт. p. 4. 22 3 дь 21 bRT. % 42°вт. -